# An Introduction on Solar Imaging Data Analytic Challenges and Opportunities for Statisticians

### v1

@@ 1.4 Challenges for Solar Flare Forecasting
The complexity of solar flares and the infrequent occurrence of energetic events results in highly heterogeneous data with large variability, making fast and accurate predictions of the time and intensity multiple hours/days ahead an extremely challenging task. According to the NOAA (2018), during solar cycle 24 (roughly 11 years), there were M flares and less than X flares. Figure 3 shows the sunspot number, which characterizes the solar activity levels, of multiple solar cycles. We can see that the different phases of the solar cycle demonstrate significantly different levels of solar activities. Our high-quality observational data to be described in the next section typically only cover one or a little more than one solar cycle.
Moreover, since A and B class flares are relatively weak and close to background levels, many of them are not recorded, causing significant “missing data”. Off-the-shelf machine learning algorithms cannot achieve desirable performance thus more delicate modeling is required to carefully extract the maximum amount of information from the complex structured data. What exacerbates the situation for data-driven methods, especially complex modeling, is the computational cost required to process the high-resolution and high cadence observations over an extended period of time. We will detail all of these aspects when introducing solar data sets in section 2.
Scientifically, it is important to (a) identify the solar active regions that have high potential to erupt in an automated fashion, (b) extract features from observed solar images using principled statistical algorithms, (c) most importantly, provide a probabilistic forecast for the eruption time, magnitude and magnetic field configuration, and (d) facilitate new understandings of the mechanism/physics of space weather events. In order to achieve these goals, we need careful statistical modeling of the evolutionary processes that finally does or does not lead to rare/extreme events; and advanced computational techniques to handle the massive, multi-faceted data sets.
@@ 2 Massive Solar Imaging Data
There is a wide spectrum of data given by various observatories with a multitude of information covered about solar activities. The data features, quality, and accessibility vary. In this section, we introduce potential available sources of data and give concrete processed data that machine learning researchers have adopted. Figure 4 shows the Sun-Earth interactions and layers of the solar environment and near-Earth environment. Figure 5 shows the NASA Heliophysics Systems Observatory that are located near-Earth or near-Sun.
@@ 2.1 Overview of Available Data Sources
As shown in Figures 4 and 5, there is a wide spectrum of observatories taking measurements about the Sun and the near-Earth environment, providing thorough information about the Sun-Earth interactions. Here, we review three major types of solar observations that are relevant for monitoring and forecasting solar activities. These observations are the ones that are most popularly adopted by scientists hoping to use data-driven approaches for space-weather monitoring. The most relevant are the imaging data, which depict solar evolution in the solar photosphere (visible “surface” of the Sun), chromosphere (the portion of solar atmosphere above the photosphere), and corona (the outer part of the solar atmosphere). Imaging data may originate from spaceborne and ground-based instruments. Most of the imaging data are taken at or near Earth and provide the Sun’s view from Sun-Earth vantage (or direction). The most notable examples include GONG, SDO, SOHO, SOLIS, STEREO (which has been on the far side of the Sun from Earth), GOES, Hinode, and Yohkoh (see, Table 1). The direct imaging data are used to (i) identify and classify solar features (sunspots and active regions, coronal holes, filaments, CMEs etc.), and (2) to create catalogs of events (e.g., solar flare catalog by NOAA, Yohkoh, and Hinode; and SOHO LASCO CME catalogs, CDAW (https://cdaw.gsfc.nasa.gov/CME_list/index.html) and Cactus (https://www.sidc.be/cactus/). GOES X-ray flux records are used to identify the X-ray flares and their properties.
@@
GONG111NSF’s Global Oscillation Network Group | Groundbased | LOS magnetograms, | 1995-P | www.gong.nso.edu |
@@
SOLIS222NSF’s Synoptic Optical Long-term Investigations of the Sun | Groundbased | Vector/LOS magnetograms, | 2003-2017 | solis.nso.edu |
@@
GONG is a global network of six solar robotic telescopes located in Australia, India, Canary Islands (Spain), Chile, and USA (California and Hawai’i). GONG provides full disk images of the solar photosphere (broadband, or white light) and chromosphere (H), line-of-sight (LOS) magnetograms and Dopplergrams (helioseimology data) in the photosphere. Helioseismology data are used to derive the so-called farside imaging (low resolution maps of solar magnetic fields on the side of the Sun opposite to one facing the Earth). The Michelson Doppler Imager (MDI) on board SOHO also produced LOS magnetograms and Doppler maps similar to GONG, albeit with a lower spatial resolution. The Extreme ultra-violet Imaging Telescope (EIT) and the Large Angle and Spectrometric Coronagraph (LASCO) are two other instruments on board SOHO spacecraft to provide images of the solar corona. The Helioseismic and Magnetic Imager (HMI) measures 3D magnetograms and the Atmospheric Imaging Assembly (AIA) photograph the Sun’s atmosphere and the corona, in multiple wavelengths.
@@ 2.2 H-alpha Images: GONG and others
From the observation perspective, full disk H (Hydrogen–) images can be used to identify the chromospheric filaments, solar flares, and their properties in the visible wavelength range. H is one of the commonly used spectral lines in solar astronomy, and thus, there is a number of ground-based instruments that provide such observations.
@@
As some examples, one can mention the Chromospheric Telescope (ChroTel, Kentischer et al., 2008; Bethge et al., 2011, years of operations 2012–2020), H instruments at the Global Oscillation Network Group (GONG, Harvey et al., 1996; Hill, 2018; NSO Integrated Synoptic Program, 2010, 2010–present), the Kanzelhöhe Observatory (KSO, Otruba and Pötzi, 2003; Pötzi et al., 2015, 2021). For additional details of these instruments and their application in machine learning, see, e.g., Diercke et al. (2024). ChroTel and H instruments at KSO are single site instruments, and thus, their observations are limited to a day-night cycle, and weather permitting. GONG is a 6-site global network, which typically provides a 90% duty cycle (90% of 24-hour period is covered by observations). Each GONG site takes H observation every 60 seconds. See Figure 6 for a snapshot of GONG H- data updated by the NSO website on May 3, 2024. The time of observations at the adjacent sites is shifted by 20 seconds, which allows to achieve a network cadence of 20 seconds. The GONG data is used by NOAA Space Weather Prediction Center (SWPC), the US Airforce 557th Weather Wing, and the NASA Community Coordinated Modeling Center (CCMC) to monitor space weather conditions.
@@
KSO H- instrument is part of the Global H Network (Steinegger et al., 2000, GHN), which is the adhoc network of telescopes operated by different organizations. The data produced by these instruments are non-uniform in both spatial resolution and spectral bands, which may lead to a difference appearance of solar features. Other notable examples are the U.S. Air Force Solar Optical Observing Network (Neidig et al., 1998, SOON), Kislovodsk High-Altitude Station of Pulkovo Observatory (KHASPO, Russia), Huairou Solar Observating Station (HSOS, P.R. China), Hida Observatory (Kyoto University), National Astronomical Observatory of Japan (NAOJ), Catania Astrophysical Observatory (CAO), Solar Survey Archive BASS2000 (SSABASS2000, France), H- telescope at Big Bear Solar Observatory (HBBSO, California). Early observations in H- were taken on the photographic film. Currently digitized (large) datasets include Solar Digitization Project at New Jersey Institute of Technology (SDP, NJIT), Kodaikanal Solar Observatory (KSO, India), NSO flare patrol telescope (NSO FPT). All groundbased observations are subject of local weather and instrument shutdown/repairs. Some digitized (scanned) imaging data are still missing digitization of meta data (e.g., time and date of observations), which could benefit tremendously from the application of pattern/text recognition and machine learning. See Table 2 for links to these data sources.
@@ 2.3 GOES: Solar Flare List and X-ray Flux
Solar flare events are recorded in the NOAA Geostationary Operational Environmental Satellites (GOES) flare list (Garcia, 1994). The flare list is popularly used by the space weather community, while some recent works show that the list miss several major flare events (Van der Sande et al., 2022). The number of solar flare events is not huge for training large-scale machine learning models with the amount of solar imaging data that we have. For example, over the period of 05/01/2010 to 06/20/2018, there are solar flares listed with class, start, end, and peak intensity time of each event. There are five levels of flare events, A/B/C/M/X, ranging from the weakest to the strongest. Table 3 gives the number of flares of the B/C/M/X class recorded in the GOES data set. The A/B are considered as weak flares, which do not impact the Earth much whereas the M/X are considered strong flare events thus early detection is crucial.
@@
Figure 7 shows the 1-minute GOES X-ray flux data that is updated in real time from SWPC/NOAA, while tracking solar flare events. The SWPC/NOAA definition of start, peak, and end time of an X-ray event is as follows. “The begin time of an X-ray event is defined as the first minute, in a sequence of 4 minutes, of steep monotonic increase in 0.1-0.8 nm flux. The X-ray event maximum is taken as the minute of the peak X-ray flux. The end time is the time when the flux level decays to a point halfway between the maximum flux and the pre-flare background level.” In the duration of 04/27/2024-05/03/2024 (in Figure 7), there is one M4.4 flare occurring.
@@ 2.4 SDO: 3D Multi-channel Solar Imaging Data
Since its launch in 2010, NASA’s Solar Dynamics Observatory (SDO; Pesnell et al. 2012) has continuously observed solar activity, providing an extensive array of scientific data for heliophysics research. There are three instruments onboard: The Atmospheric Imaging Assembly (AIA; Lemen et al. 2012) captures full-disk images of the Sun in high spatial (40964096, pixel size of 0.6 arcseconds) and high temporal (12 seconds for EUV channels 94, 131, 171, 193, 211, 304, and 335 Å, and 24 seconds for UV channels 1600 and 1700 Å) resolution. The Helioseismic and Magnetic Imager (HMI; Scherrer et al. 2012) captures visible wavelength filtergrams of the full Sun at 40964096 resolution (pixel size of 0.5 arcsec). These filtergrams are then processed into various data products, including photospheric Dopplergrams, line-of-sight magnetograms, and vector magnetograms (Hoeksema et al., 2014). The EUV Variability Experiment (EVE; Woods et al. 2012) measures the solar EUV spectral irradiance from 1 to 1050 Å. Figure 8 shows an example of the HMI and AIA full disc data. Tables 4 and 5 give more information on the coverage and properties of the SDO data.
@@
1600 Å | 200 | C iv, continumm | Transition region, upper photosphere | 5.0 |
@@
While SDO data are easily accessible, pre-processing these data for a scientific analysis often requires specialized heliophysics knowledge. To facilitate the SDO data usage, Galvez et al. (2019) created a curated data set from the SDO mission in a format suitable for machine learning research. The SDOML Dataset is preprocessed from the original Level 1 data by down-sampling AIA and HMI images from 40964096 to 512512 pixels, removed observations, corrected for instrumental degradation over time, and applied exposure corrections. Both AIA and HMI data are spatially colocated, have identical angular resolutions, and that all instruments are chronosynchronous. The temporal resolution for AIA is 6 minutes. The temporal resolution for HMI vector magnetic field observations in Bx, By, and Bz components is 12 minutes. The EVE observations in 39 wavelengths from 2010-05-01 to 2014-05-26. The temporal resolution is 10 second. The dataset is permanently stored at Stanford Digital Repository in .npz format. There is a recent update on this dataset, in which the full dataset has been converted to cloud friendly Zarr format with full FITS hearder information. Both v1 and v2 version of the dataset is now available through NASA HelioCloud on AWS (https://registry.opendata.aws/sdoml-fdl/).
@@
Previous work has established that solar eruptions are all associated with highly nonpotential magnetic fields that store the necessary free energy. The most energetic flares come from very localized intense kilogauss fields of Active Regions (ARs) Forbes (2000); Schrijver (2009). At any time, these ARs occupy a small area of the solar surface. The magnetic complexity of active regions and its evolution are often serve as the precursors of their flaring potential. Many CMEs are associated with the activation and subsequent eruption of the chromospheric filaments. Thus, the current applications of machine learning in solar and heliospheric physics are concentrating on identifying these solar phenomena. The full disk HMI magnetogram data are subdivided into much smaller AR patches, where several of them often occur simultaneously, see Figure 9 for the active region patches. Therefore, when handling AR-only data, the number of subdivided images that we have is another order of magnitude larger whereas the amount of total data is significantly reduced (from 6 TB to 413 GB).
From each SHARP patch, scalar parameters (called SHARP parameters) are calculated to capture the zeroth order structure and complexity of the magnetic field. As discussed in Leka and Barnes (2003) and Bobra et al. (2014a), these parameters are designed to assess the flaring potential of active regions and are thus strongly representative of the total free energy of the magnetic field. These whole-active-region magnetic quantities can be effectively used as predictors of flares and also CMEs (cf. Falconer, 2001; Falconer et al., 2002b, 2003, 2006; Leka and Barnes, 2003; Schrijver, 2007; Bobra and Couvidat, 2015).
@@
In our data pre-processing pipeline (Chen et al., 2019), the flare events are matched to the Space-weather HMI Active Region Patch (SHARP) vector field data patches provided by the Joint Science Operations Center (JSOC) website. The SHARP patches contain 2-D photospheric maps of orthogonal magnetic field components observed by the Helioseismic and Magnetic Imager (HMI) on Solar Dynamics Observatory (SDO) with 1.0 arcsecond spatial resolution (4096x4096 pixel images) and time cadence of 12 minutes (Hoeksema et al., 2014; Bobra et al., 2014a). This gives us high-resolution images per day from the three channels, which amounts to million high resolution images over the year period that we work on. Due to the rotation of the Sun, an active region cannot be seen clearly (within 68 degrees of central meridian to avoid foreshortening) for more than approximately 250 hours, which corresponds to time frames, at a time.

### v2

@@ 1.4 Challenges for Solar Flare Forecasting
The complexity of solar flares and the infrequent occurrence of energetic events results in highly heterogeneous data with large variability, making fast and accurate predictions of the time and intensity multiple hours/days ahead an extremely challenging task. According to the NOAA (2018), during solar cycle 24 (roughly 11 years), there were M flares and less than X flares. Figure 3 shows the sunspot number, which characterizes the solar activity levels of multiple solar cycles. We can see that the different phases of the solar cycle demonstrate significantly different levels of solar activities. Our high-quality observational data to be described in the next section typically only cover one or a little more than one solar cycle.
Moreover, since A and B class flares are relatively weak and close to background levels, many of them are not recorded, causing significant “missing data”. Off-the-shelf machine learning algorithms cannot achieve desirable performance; thus, more delicate modeling is required to carefully extract the maximum amount of information from the complex structured data. What exacerbates the situation for data-driven methods, especially complex modeling, is the computational cost required to process the high-resolution and high-cadence observations over an extended period. We will detail these aspects when introducing solar data sets in section 2.
Scientifically, it is important to (a) identify the solar active regions that have high potential to erupt in an automated fashion, (b) extract features from observed solar images using principled statistical algorithms, (c) most importantly, provide a probabilistic forecast for the eruption time, magnitude and magnetic field configuration, and (d) facilitate new understandings of the mechanism/physics of space weather events. To achieve these goals, we need (1) careful statistical modeling of the spatial and temporal patterns, thus identifying precursors that finally do (or do not) lead to rare and extreme events, and (2) advanced computational techniques to handle the massive and multi-faceted data sets.
@@ 2 Massive Solar Imaging Data
There is a broad spectrum of data given by various observatories, and a multitude of information is covered about solar activities. Figure 4 shows the Sun-Earth interactions and layers of the solar and near-Earth environments. Figure 5 shows the NASA Heliophysics Systems Observatory located near-Earth or near-Sun. The data features, quality, and accessibility vary across observatories. In this paper, we focus on solar imaging data, which are typically given by the Flexible Image Transport System (FITS), i.e., as multidimensional numerical arrays. This section introduces potential available data sources and gives concrete processed data that machine learning researchers have adopted.
@@ 2.1 Overview of Available Data Sources
As shown in Figures 4 and 5, there is a broad spectrum of observatories taking measurements about the Sun and the near-Earth environment, providing thorough information about the Sun-Earth interactions. Here, we review three major types of solar observations relevant to monitoring and forecasting solar activities. These observations are the ones most popularly adopted by scientists hoping to use data-driven approaches for space-weather monitoring. The most relevant are the imaging data, which depict solar evolution in the solar photosphere (visible “surface” of the Sun), chromosphere (the portion of solar atmosphere above the photosphere), and corona (the outer part of the solar atmosphere). Imaging data may originate from spaceborne and ground-based instruments. Most imaging data are taken at or near Earth and provide the Sun’s view from Sun-Earth vantage (or direction). The most notable examples include GONG, SDO, SOHO, SOLIS, STEREO (which has been on the far side of the Sun from Earth), GOES, Hinode, and Yohkoh (see, Table 1). The direct imaging data are used to (i) identify and classify solar features (sunspots and active regions, coronal holes, filaments, CMEs, etc.) and (2) to create catalogs of events (e.g., solar flare catalog by NOAA, Yohkoh, and Hinode; and SOHO LASCO CME catalogs, CDAW (https://cdaw.gsfc.nasa.gov/CME_list/index.html) and Cactus (https://www.sidc.be/cactus/). GOES X-ray flux records are used to identify the X-ray flares and their properties.
@@
GONG111NSF’s Global Oscillation Network Group | Ground-based | LOS magnetograms, | 1995-P | www.gong.nso.edu |
@@
SOLIS222NSF’s Synoptic Optical Long-term Investigations of the Sun | Ground-based | Vector/LOS magnetograms, | 2003-2017 | solis.nso.edu |
@@
GONG is a global network of six solar robotic telescopes located in Australia, India, the Canary Islands (Spain), Chile, and the USA (California and Hawai’i). GONG provides full disk images of the solar photosphere (broadband or white light) and chromosphere (H), line-of-sight (LOS) magnetograms, and Dopplergrams (helioseismology data) in the photosphere. Helioseismology data are used to derive the so-called farside imaging (low-resolution maps of solar magnetic fields on the side of the Sun opposite to one facing the Earth). The Michelson Doppler Imager (MDI) on board SOHO also produced LOS magnetograms and Doppler maps similar to GONG, albeit with a lower spatial resolution. The Extreme Ultra-violet Imaging Telescope (EIT) and the Large Angle and Spectrometric Coronagraph (LASCO) are other instruments on board SOHO spacecraft providing solar corona images. The Helioseismic and Magnetic Imager (HMI) measures 3D magnetograms, and the Atmospheric Imaging Assembly (AIA) photographs the Sun’s atmosphere and the corona in multiple wavelengths.
@@ 2.2 H-alpha Images: GONG and others
From the observation perspective, full disk H (Hydrogen–) images can be used to identify the chromospheric filaments, solar flares, and their properties in the visible wavelength range. H is one of the commonly used spectral lines in solar astronomy, and thus, there are a number of ground-based instruments that provide such observations.
@@
As some examples, one can mention the Chromospheric Telescope (ChroTel, Kentischer et al., 2008; Bethge et al., 2011, years of operations 2012–2020), H instruments at the Global Oscillation Network Group (GONG, Harvey et al., 1996; Hill, 2018; NSO Integrated Synoptic Program, 2010, 2010–present), the Kanzelhöhe Observatory (KSO, Otruba and Pötzi, 2003; Pötzi et al., 2015, 2021). For additional details of these instruments and their application in machine learning, see, e.g., Diercke et al. (2024). ChroTel and H instruments at KSO are single-site instruments, and thus, their observations are limited to a day-night cycle, weather permitting. GONG is a 6-site global network, which typically provides a 90% duty cycle (90% of the 24-hour period is covered by observations). Each GONG site takes H observation every 60 seconds. See Figure 6 for a snapshot of GONG H- data updated by the NSO website on May 3, 2024. The time of observations at the adjacent sites is shifted by 20 seconds, which allows for the achievement of a network cadence of 20 seconds. The GONG data is used by the NOAA Space Weather Prediction Center (SWPC), the US Air Force 557th Weather Wing, and the NASA Community Coordinated Modeling Center (CCMC) to monitor space weather conditions.
@@
KSO H- instrument is part of the Global H Network (Steinegger et al., 2000, GHN), which is the adhoc network of telescopes operated by different organizations. The data produced by these instruments are non-uniform in both spatial resolution and spectral bands, which may lead to a difference in the appearance of solar features. Other notable examples are the U.S. Air Force Solar Optical Observing Network (Neidig et al., 1998, SOON), Kislovodsk High-Altitude Station of Pulkovo Observatory (KHASPO, Russia), Huairou Solar Observing Station (HSOS, P.R. China), Hida Observatory (Kyoto University), National Astronomical Observatory of Japan (NAOJ), Catania Astrophysical Observatory (CAO), Solar Survey Archive BASS2000 (SSABASS2000, France), H- telescope at Big Bear Solar Observatory (HBBSO, California). Early observations in H- were taken on the photographic film. Currently, digitized (large) datasets include the Solar Digitization Project at New Jersey Institute of Technology (SDP, NJIT), Kodaikanal Solar Observatory (KSO, India), and NSO flare patrol telescope (NSO FPT). All ground-based observations are subject to local weather and instrument shutdown/repairs. Some digitized (scanned) imaging data are still missing digitization of metadata (e.g., time and date of observations), which could benefit tremendously from the application of pattern/text recognition and machine learning. See Table 2 for links to these data sources.
@@ 2.3 GOES: Solar Flare List and X-ray Flux
Solar flare events are recorded in the NOAA Geostationary Operational Environmental Satellites (GOES) flare list (Garcia, 1994). The flare list is popularly used by the space weather community, while some recent works show that the list misses several major flare events (Van der Sande et al., 2022). The number of solar flare events is not massive enough to train large-scale machine-learning models with the amount of solar imaging data we have. For example, from 05/01/2010 to 06/20/2018, solar flares are listed with class, start, end, and peak intensity time of each event. There are five levels of flare events, A/B/C/M/X, ranging from the weakest to the strongest. Table 3 gives the number of flares of the B/C/M/X class recorded in the GOES data set. The A/B are considered weak flares, which do not impact the Earth much, whereas the M/X are considered strong flare events; thus, early detection is crucial.
@@
Figure 7 shows the 1-minute GOES X-ray flux data updated in real-time from SWPC/NOAA while tracking solar flare events. The SWPC/NOAA definition of the start, peak, and end time of an X-ray event is as follows. “The begin time of an X-ray event is defined as the first minute, in a sequence of 4 minutes, of steep monotonic increase in 0.1-0.8 nm flux. The X-ray event maximum is taken as the minute of the peak X-ray flux. The end time is when the flux level decays to a point halfway between the maximum flux and the pre-flare background level.” In the duration of 04/27/2024-05/03/2024 (in Figure 7), there is one M4.4 flare occurring.
@@ 2.4 SDO: 3D Multi-channel Solar Imaging Data
Since its launch in 2010, NASA’s Solar Dynamics Observatory (SDO; Pesnell et al. 2012) has continuously observed solar activity, providing an extensive array of scientific data for heliophysics research. There are three instruments onboard: The Atmospheric Imaging Assembly (AIA; Lemen et al. 2012) captures full-disk images of the Sun in high spatial (40964096, pixel size of 0.6 arcseconds) and high temporal (12 seconds for EUV channels 94, 131, 171, 193, 211, 304, and 335 Å, and 24 seconds for UV channels 1600 and 1700 Å) resolution. The Helioseismic and Magnetic Imager (HMI; Scherrer et al. 2012) captures visible wavelength filtergrams of the full Sun at 40964096 resolution (pixel size of 0.5 arcsec). Note that AIA has a slightly larger FOV (Field of View) (41 arcmin) than HMI (34 arcmin), which leads to the different pixel sizes mentioned above. These filtergrams are then processed into various data products, including photospheric Dopplergrams, line-of-sight magnetograms, and vector magnetograms (Hoeksema et al., 2014). The EUV Variability Experiment (EVE; Woods et al. 2012) measures the solar EUV spectral irradiance from 1 to 1050 Å. Figure 8 shows an example of the HMI and AIA full disc data. Tables 4 and 5 give more information on the coverage and properties of the SDO data.
@@
1600 Å | 200 | C iv, continuum | Transition region, upper photosphere | 5.0 |
@@
While SDO data are easily accessible, pre-processing these data for a scientific analysis often requires specialized heliophysics knowledge. To facilitate the SDO data usage, Galvez et al. (2019) created a curated data set from the SDO mission in a format suitable for machine learning research. The SDOML Dataset is preprocessed from the original Level 1 data by down-sample HMI and AIA images from 40964096 to 512512 pixels, removed observations, corrected for instrumental degradation over time, and applied exposure corrections. Both AIA and HMI data are spatially colocated and have identical angular resolutions (pixel size of 4.8 arcsec) , and that all instruments are chronosynchronous. The temporal resolution for AIA is 6 minutes. The temporal resolution for HMI vector magnetic field observations in Bx, By, and Bz components is 12 minutes. The EVE observations in 39 wavelengths from 2010-05-01 to 2014-05-26. The temporal resolution is 10 seconds. The dataset is permanently stored at the Stanford Digital Repository in .npz format. There is a recent update on this dataset, in which the entire dataset has been converted to a cloud-friendly Zarr format with complete FITS header information. Both the v1 and v2 versions of the dataset are now available through NASA HelioCloud on AWS (https://registry.opendata.aws/sdoml-fdl/).
@@
Previous work has established that solar eruptions are all associated with highly nonpotential magnetic fields that store the necessary free energy. The most energetic flares come from very localized intense kilogauss fields of Active Regions (ARs) Forbes (2000); Schrijver (2009). At any time, these ARs occupy a small area of the solar surface. The magnetic complexity of active regions and their evolution often serve as the precursors of their flaring potential. Many CMEs are associated with the activation and subsequent eruption of the chromospheric filaments. Thus, the current applications of machine learning in solar and heliospheric physics are concentrating on identifying these solar phenomena. The full disk HMI magnetogram data are subdivided into much smaller AR patches, where several of them often occur simultaneously; see Figure 9 for the active region patches. Therefore, when handling AR-only data, the number of subdivided images that we have is another order of magnitude larger. In contrast, the amount of total data is significantly reduced (from 6 TB to 413 GB).
From each SHARP patch, scalar parameters (called SHARP parameters) are calculated to capture the zeroth order structure and complexity of the magnetic field. As discussed in Leka and Barnes (2003) and Bobra et al. (2014a), these parameters are designed to assess the flaring potential of active regions. They are thus strongly representative of the total free energy of the magnetic field. These whole-active-region magnetic quantities can be effectively used as predictors of flares and also CMEs (cf. Falconer, 2001; Falconer et al., 2002b, 2003, 2006; Leka and Barnes, 2003; Schrijver, 2007; Bobra and Couvidat, 2015).
@@
In our data pre-processing pipeline (Chen et al., 2019), the GOES flare list is matched to the Space-weather HMI Active Region Patch (SHARP) vector field data patches provided by the Joint Science Operations Center (JSOC) website. The SHARP patches contain 2-D photospheric maps of orthogonal magnetic field components observed by the Helioseismic and Magnetic Imager (HMI) on Solar Dynamics Observatory (SDO) with 1.0 arcsecond spatial resolution (4096x4096 pixel images) and time cadence of 12 minutes (Hoeksema et al., 2014; Bobra et al., 2014a). This gives us high-resolution images per day from the three channels, which amounts to million high resolution images over the year period that we work on. Due to the rotation of the Sun, an active region cannot be seen clearly (within 68 degrees of the central meridian to avoid foreshortening) for more than approximately 250 hours, which corresponds to time frames at a time.